OpenStudy (anonymous):

The number of plant species listed as threatened (likely to become endangered) or endangered (in danger of becoming extinct) has more than tripled in past 20 years. In 2008, there were 746 species of plants in the United States that were considered threatened or endangered. The number considered threatened was 4 less than one-fourth of the number considered endangered. How many U.S. plant species were considered endangered and how many were considered threatened in 2008

OpenStudy (immings):

If you are looking for an equation it is a little bit different because of the 746 this is considered threatened OR endangered, not just endangered. Therefore we cannot give a specific number E=endangered T=Threatened T=[0.25xE]-4 If you use (746x0.25)-4=182.5 or 183 I don't think this would be the accurate number because the 746 is for both. sorry I'm not 100% that this (183) correct.

OpenStudy (anonymous):

This is what I got, but not sure what is next. t+d=746 t=(1/4)d-4